The outputs of four registers, A, B, C, D, H, L are connected through 8:1 line
multiplexer to the input of a fifth register, X. Each register is 16 bits long. The
required transfers are dictated by four timing variables, T0 through T3 as follows:
T0 : X←A
T1 : X←B
T2 : X←C
1
Expert's answer
2021-09-23T03:05:30-0400

Given:

T0 : X←A

T1 : X←B

T2 : X←C

The truth table for the outputs:



Use the “sum of product” technique to get the logic expressions for the selection lines and the load signal: S1 = T2 + T3

S0 = T1 + T3

Load = T0 + T1 + T2 + T3

From this, we can illustrate as follow: